OpenStudy (anonymous):

How do you find the two points from this equation to be graphed? 3y+6=0

OpenStudy (anonymous):

do you mean 3x ? or is it actually 3y?

OpenStudy (anonymous):

its 3y

OpenStudy (rock_mit182):

if the equation is \[ 3y+6= 0\] then

OpenStudy (rock_mit182):



OpenStudy (anonymous):

so would you write this as (-1,1), (-1,2) ?

OpenStudy (rock_mit182):

the only value that make the equation true is when Y equals - 2, since: 3(-2) +6 =0 -6+6 =0 0 = 0 which means the function is constant, in other words Y equals -2 no matter what